• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name eschulte

    Log entry time 14:23:29 on February 26, 2011

    Entry number 348827

    keyword=Still in the Hall

    We have received no update in the counting house as to the status of the spectrometer motion. I'm assuming we're still working on it, since they haven't moved.

    Also, "the guys" are working on the dead-time issues we've been seeing.